ADNAN MUHAMMED

Grade 12,

If alpha and beta are the zeroes of a quadratic polynomial 3x^2+5x+7,then 1/alpha cube + 1/b cube is?

If alpha and beta are the zeroes of a quadratic polynomial 3x^2+5x+7,then 1/alpha cube + 1/b cube is?

Grade:10

1 Answers

Harsh Patodia IIT Roorkee
6 years ago
here alpha+beta=-5/3
alpha*beta=7/3
now 1/a3+1/b3=a3+b3/a3b3=(a+b)(a2-ab+b2)/a3b3=(a+b){(a+b)2-3ab}/a3b3
=(-5/3){(-5/3)2-3(7/3)}/(7/3)3=-20/343